Calculate the number of atoms in a 3.03 x 10^3 g sample of carbon.
1
Expert's answer
2019-09-20T07:36:02-0400

m = 3.03 x 10^3 g = 3030 g

n(C) = m/M = 3030 g /(12 g/mol) = 252.5 mol